# Manifold-aware Representation Learning for Degradation-agnostic Image Restoration

### v1

@@ Abstract
Image Restoration (IR) aims to recover high-quality images from degraded inputs affected by various corruptions such as noise, blur, haze, rain, and low-light conditions. Despite recent advances, most existing approaches treat IR as a direct mapping problem, relying on shared representations across degradation types without modeling their structural diversity. In this work, we present MIRAGE, a unified and lightweight framework for all-in-one IR that explicitly decomposes the input feature space into three semantically aligned parallel branches, each processed by a specialized module—attention for global context, convolution for local textures, and MLP for channel-wise statistics. This modular decomposition significantly improves generalization and efficiency across diverse degradations. Furthermore, we introduce a cross-layer contrastive learning scheme that aligns shallow and latent features to enhance the discriminability of shared representations. To better capture the underlying geometry of feature representations, we perform contrastive learning in a Symmetric Positive Definite (SPD) manifold space rather than the conventional Euclidean space. Extensive experiments show that MIRAGE not only achieves new state-of-the-art performance across a variety of degradation types but also offers a scalable solution for real-world IR scenarios. Our code and models will be publicly available at our Project Page.
@@ 1 Introduction
Image Restoration (IR) is a fundamental yet challenging problem in computer vision, aiming to reconstruct clean images from their degraded versions affected by a variety of real-world corruptions, including noise, blur, haze, rain, low-light conditions, and more zamir2022restormer ; li2023efficient ; ren2024sharing ; zamfir2025moce ; ma2024shapesplat ; potlapalli2023promptir . While recent advances have made great strides by designing powerful architectures based on Convolutional Neural Networks (CNNs), Multi-Layer Perceptrons (MLPs), and Transformers dosovitskiy2020image ; ren2024sharing ; zhu2023good , most existing methods view restoration as a direct mapping problem—learning a global function that transforms the corrupted input into its clean counterpart. These approaches typically rely on shared representations across all degradation types, often overlooking the structural distinctions among them. As a result, they struggle to generalize when confronted with unseen or mixed degradations.
In reality, different degradation types, such as noise, haze, rain, snow, and motion blur, arise from fundamentally different physical processes and can be roughly categorized as additive, multiplicative, or kernel-based corruptions. Simultaneously, basic architectural modules exhibit distinct processing biases: convolutional filters excel at modeling local texture, attention mechanisms capture global dependencies, and MLPs often enhance channel-wise statistics. This leads to a key insight: To handle diverse and complex degradations effectively, a restoration model should be equipped with multiple processing capabilities. However, designing such a unified framework is non-trivial. A naive combination of CNNs, MLPs, and attention units often leads to redundancy, optimization difficulty, and computational inefficiency. In this work, we revisit the widely observed redundancy in attention mechanisms, particularly along the channel dimension venkataramanan2023skip ; ji2025always ; dong2021attention . This phenomenon has been extensively validated in both the natural language processing (NLP) and vision communities, showing that many attention heads or channels contribute minimally to performance. Rather than avoiding this redundancy, we propose to exploit it to build an efficient, modular, and generalizable architecture. We present MIRAGE, a unified and lightweight IR framework that explicitly decomposes feature processing into three semantically aligned branches. Specifically, we split the input feature map along the channel dimension and assign each partition to a specialized module: attention for global context modeling, convolution for local texture enhancement, and MLP for channel-wise refinement. This structured decomposition not only enhances expressiveness but also reduces computational complexity, yielding a powerful backbone that performs competitively across a wide range of degradation scenarios.
While MIRAGE already provides strong results and surpasses many state-of-the-art methods, we observe that its performance primarily stems from its architectural expressiveness rather than an explicit understanding of degradation-specific semantics. Motivated by recent efforts such as AirNet and PromptIR, which incorporate degradation cues into the learning process, we argue that degradation-guided processing can make the shared representation more discriminative and significantly improve generalization across diverse restoration tasks. To this end, we introduce a novel cross-layer contrastive learning strategy. Unlike conventional contrastive learning, which relies on complex augmentations, heavy architectures, or multi-stage training, we take a simple yet effective approach. Inspired by deeply supervised networks lee2015deeply , we hypothesize that natural contrastive pairs exist within the model itself, i.e., between shallow and deeper latent representations. These layers carry complementary characteristics: shallow features are rich in spatial details and noise sensitivity, while latent features are more abstract and semantically consistent. Aligning them can improve both robustness and generalization. Furthermore, instead of applying contrastive loss in Euclidean space, which can misrepresent feature similarity, we perform it in the Symmetric Positive Definite (SPD) manifold space, leading to more meaningful geometric alignment.
Inspired by the metaphor of uncovering the latent clean scene beneath complex degradations, just as a mirage reveals a hidden reality, our method learns a degradation-agnostic representation by dynamically balancing global context, local structure, and channel-wise distribution. This synergy between architectural modularity and contrastive regularization makes MIRAGE a powerful and general-purpose backbone for challenging all-in-one image restoration.
Our main contributions are as follows:
  • We propose MIRAGE, a lightweight and unified IR framework that synergistically integrates attention, convolution, and MLP through structured channel-wise decomposition, enabling robust performance across diverse degradation types.
  • We introduce a simple yet effective cross-layer contrastive learning strategy that aligns shallow and latent features within the SPD manifold space, thereby enhancing feature discriminability and generalization. Our approach respects the intrinsic geometry of deep representations and improves restoration quality, all without relying on data augmentation or multi-stage training.
  • Extensive experiments across diverse restoration tasks demonstrate the effectiveness and efficiency of our approach. We hope MIRAGE can serve as a strong and practical baseline to inspire future research in all-in-one image restoration.
@@ 2 Related Work
Image Restoration (IR) with Various Basic Architectures. IR aims to address a highly ill-posed problem: reconstructing high-quality images from their degraded versions. Given its broad importance, IR has been widely applied in numerous applications richardson1972bayesian ; xie2025mat ; banham1997digital ; li2023lsdir ; zamfir2024details . Early IR methods primarily relied on model-based solutions that searched for closed-form solutions to predefined formulations. With the advent of deep neural networks, learning-based approaches have rapidly gained popularity. A wide range of methods have emerged, including regression-based techniques lim2017enhanced ; lai2017deep ; liang2021swinir ; chen2021learning ; li2023efficient ; zhang2024transcending and generative model-based pipelines gao2023implicit ; wang2022zero ; luo2023image ; yue2023resshift ; zhao2024denoising , built upon convolutional dong2015compression ; zhang2017learning ; zhang2017beyond ; wang2018recovering , MLP-based tu2022maxim , state space models guo2024mambair ; zhu2024vision ; mamba ; mamba2 , and Vision Transformer (ViT)-based architectures liang2021swinir ; ren2023masked ; li2023efficient ; zamir2022restormer ; dosovitskiy2020image ; liu2023spatio . Despite the impressive performance of recent state-of-the-art methods, most IR solutions are still designed to address specific degradation types, such as denoising zhang2017learning ; zhang2019residual , dehazing ren2020single ; wu2021contrastive , deraining jiang2020multi ; ren2019progressive , deblurring kong2023efficient ; ren2023multiscale , and others.
@@
Degradation-agnostic Image Restoration. While training task-specific models to handle individual types of degradation can be effective, it poses practical limitations due to the need for separate models for each degradation. In real-world scenarios, images often suffer from a mixture of degradations and artifacts, making it difficult to address each type independently. Task-specific solutions also demand considerable computational and storage resources, significantly increasing their environmental footprint. To overcome these limitations, the emerging field of All-in-One image restoration focuses on single-blind models capable of handling multiple degradation types simultaneously zamfir2025moce ; zeng2025vision ; zheng2024learning . For example, AirNet li2022all achieves blind All-in-One image restoration by using contrastive learning to derive degradation representations from corrupted images, which are then leveraged to reconstruct clean images. Building on this, IDR zhang2023ingredient tackles the problem by decomposing degradations into fundamental physical components and applying a two-stage meta-learning strategy. More recently, the prompt-based paradigm potlapalli2023promptir ; wang2023promptrestorer ; li2023prompt has introduced a visual prompt learning module, enabling a single model to better handle diverse degradation types by leveraging the discriminative capacity of learned visual prompts. Extending this idea, some works further model prompts from a frequency perspective cui2025adair or propose more complex architectures with additional datasets dudhane2024dynamic . However, visual prompt modules often result in increased training time and decreased efficiency cui2025adair . In contrast, our work aims to improve the model’s ability to capture representative degradation cues without relying on heavy or complex prompt designs. Our goal is to develop an All-in-One image restorer that remains both computationally efficient and environmentally sustainable.
3 Preliminary: Degradation-Aware Architectures for Image Restoration
@@
Image Degradation and Restoration. Image restoration aims to recover a clean image $\mathbf{x}$ from a degraded observation $\mathbf{y}$, commonly modeled as:
@@
where $\mathcal{D}(\cdot)$ is a degradation operator and $\mathbf{n}$ is noise. Real-world degradations are diverse in nature—ranging from additive (e.g., Gaussian noise, rain: $\mathbf{y}=\mathbf{x}+\mathbf{n}$), multiplicative (e.g., haze, speckle: $\mathbf{y}=\mathbf{x}\cdot\mathbf{m}$), to convolutional degradations (e.g., blur, super-resolution: $\mathbf{y}=\mathbf{k}*\mathbf{x}+\mathbf{n}$).
Such degradations are often entangled and spatially variant, forming compound pipelines:
$\mathbf{y}=\mathcal{D}_{3}\big{(}\mathcal{D}_{2}(\mathcal{D}_{1}(\mathbf{x}))\big{)}+\mathbf{n},$ (2)
where each $\mathcal{D}_{i}$ captures a distinct degradation type. Addressing this diversity requires models capable of both local detail preservation and global structural reasoning.
Architectural Biases for Degradation Modeling. To this end, modern deep networks adopt different inductive biases: CNNs capture local spatial patterns through convolution: $\mathbf{y}_{p}=\sum_{i\in\mathcal{N}(p)}w_{i}\cdot\mathbf{x}_{i}$. Their locality makes them effective for uniform or spatially invariant degradations. Transformers exploit global self-attention: $\mathbf{y}_{i}=\sum_{j}\alpha_{ij}\cdot\mathbf{V}_{j}$, enabling modeling of non-uniform, structured degradations such as haze or patterned noise. MLPs, especially token-mixing variants, use position-wise transformations: $\mathbf{y}=\mathbf{W}_{2}\cdot\phi(\mathbf{W}_{1}\cdot\mathbf{x})$, offering flexibility with weak spatial priors.
Each paradigm exhibits strengths and limitations in handling different degradation types. CNNs excel in local fidelity, Transformers in global context, and MLPs offer flexible feature interactions but lack inherent spatial priors. When applied to vision tasks, they often require large parameter counts and are less effective alone for structured degradations. These complementary traits motivate unified architectures that integrate them for robust, degradation-aware restoration in the wild.
@@ 4 The Proposed MIRAGE
Figure 2: (a)-(c): The most adopted all-in-one image restoration encodedr-decoder pipelines. (d): The toy illustration of our SPD contrastive pipeline. (e): The overall framework of the proposed MIRAGE : i.e., a convolutional patch embedding, a U-shape encoder-decoder main body, an extra refined block, and the proposed SPD contrastive learning algorithm. (f): Structure of each mixed degradation adaptation block (MDAB).
Overview. Unlike previous methods that adopt a one-to-one mapping strategy for multiple degradations as illustrated in Fig. 2(a), which requires training a separate model for each type of degradation, our approach is more unified. Also, unlike architectures that use multiple encoders mapping to a single decoder, as shown in Fig. 2(b), which significantly increases the model size and complexity, we pursue a more efficient design. Moreover, in contrast to prompt-based methods that rely on large-scale models or introduce visual/textual prompts (Fig. 2(c)), we propose a simple yet effective mixed-backbone architecture (Fig. 2(d)). This backbone serves as a strong restoration model (Sec. 4.1), and its performance is further enhanced through cross-layer contrastive learning in the SPD space between shallow and latent features (Sec. 4.2).
4.1 Mixed Degradation Adaptation Block for AnyIR
Redundancy in MHAs Opens Opportunities for Hybrid Architectures. Redundancy has long been recognized as a fundamental limitation in multi-head self-attention (MHA), the core building block of Transformers, in both NLP and vision domains nguyen2022improving ; nguyen2022improvingkeys ; xiao2024improving ; brodermann2025cafuser ; wang2022improved ; venkataramanan2023skip . Prior studies have shown that not all attention heads contribute equally, i.e., some are specialized and crucial, while others can be pruned with negligible impact. This inherently implies redundancy in the channel dimension, as MHA outputs are concatenated along this axis. To empirically verify this redundancy in the context of image restoration (IR), we analyze intermediate features from a lightweight attention-only model (details in the Appendix).
@@ 4.1 Mixed Degradation Adaptation Block for AnyIR
Specifically, we compute the cumulative explained variance via PCA and the normalized singular value spectra via SVD across multiple feature scales, as shown in Fig.3. As illustrated in Fig. 3(a), earlier scales (e.g., 1st Scale) require significantly fewer principal components to retain most of the variance, suggesting high redundancy. Fig. 3(b) further supports this observation, with a sharper singular value decay at shallower stages, indicating stronger low-rank structure in channel-wise representations. Even at the deepest stage (e.g., 4th Scale), achieving 90% variance only requires 28 out of 192 components ($\approx 19\%$), confirming that redundancy persists throughout.
This insight motivates a departure from traditional head/channel pruning. Instead of discarding redundant capacity, we propose to repurpose it by splitting the channel dimension into three parts and feeding them into distinct architectural branches, e.g., attention, convolution, and MLP. This hybrid formulation leverages complementary inductive biases and makes full use of available representational space, offering a principled and efficient alternative to the previous pure MSA-based designs.
@@
1:$F_{\text{in}}^{\text{att}},\ F_{\text{in}}^{\text{conv}},\ F_{\text{in}}^{\text{mlp}}$triangleright Input features from three branches
2:$F_{\text{out}}$triangleright Final fused output
@@
4:$Q,K,V\leftarrow\text{Linear}(F_{\text{in}}^{\text{att}})$triangleright Projection to attention tokens
5:$F_{\text{out}}^{\text{att}}\leftarrow\texttt{Softmax}(\frac{QK^{\top}}{\sqrt{d}})V$triangleright Multi-head self-attention
@@
7:$F^{\prime}\leftarrow\text{Conv1x1}(\text{Norm}(F_{\text{in}}^{\text{conv}}))$triangleright Normalization and expansion
8:$\gamma,\beta,\alpha\leftarrow\text{Split}(F^{\prime})$triangleright Gating, intermediate, convolutional paths
9:$\alpha^{\prime}\leftarrow\text{DynamicDepthwiseConv}(\alpha)$triangleright Content-adaptive depthwise conv
10:$\hat{F}\leftarrow\sigma(\gamma/\tau)\cdot\text{Concat}(\beta,\alpha^{\prime})$triangleright Gated local enhancement
11:$F_{\text{out}}^{\text{conv}}\leftarrow\text{Conv1x1}(\hat{F})+F_{\text{in}}^{\text{conv}}$triangleright Residual projection
@@
13:$F_{\text{out}}^{\text{mlp}}\leftarrow\text{MLP}(F_{\text{in}}^{\text{mlp}})$triangleright Channel-wise transformation brings more non-linearity
@@
15:$F_{\text{out}}^{\text{att}^{\prime}}\leftarrow F_{\text{out}}^{\text{att}}+\lambda_{\text{att}}\cdot\sigma(F_{\text{out}}^{\text{conv}}+F_{\text{out}}^{\text{mlp}})$triangleright Fuse conv and MLP into attention
16:$F_{\text{out}}^{\text{conv}^{\prime}}\leftarrow F_{\text{out}}^{\text{conv}}+\lambda_{\text{conv}}\cdot\sigma(F_{\text{out}}^{\text{att}}+F_{\text{out}}^{\text{mlp}})$triangleright Fuse attention and MLP into conv
17:$F_{\text{out}}^{\text{mlp}^{\prime}}\leftarrow F_{\text{out}}^{\text{mlp}}+\lambda_{\text{mlp}}\cdot\sigma(F_{\text{out}}^{\text{att}}+F_{\text{out}}^{\text{conv}})$triangleright Fuse attention and conv into MLP
@@
19:$F_{\text{out}}^{\text{fuse}}\leftarrow\text{Project}(\text{Concat}(F_{\text{out}}^{\text{att}^{\prime}},\ F_{\text{out}}^{\text{conv}^{\prime}},\ F_{\text{out}}^{\text{mlp}^{\prime}}))$triangleright Final unified representation
@@
Parallel Design Brings More Efficiency. The preceding redundancy analysis reveals that many channels across scales encode overlapping information. Rather than discarding this capacity via pruning, we strategically re-purpose it through a structurally parallel design. As illustrated in Fig. 2(f), the input feature $F_{\text{in}}\in\mathbb{R}^{h\times w\times c}$ is evenly divided along the channel dimension into three sub-tensors: $F_{\text{in}}^{\text{att}}$, $F_{\text{in}}^{\text{conv}}$, and $F_{\text{in}}^{\text{mlp}}$, corresponding to attention, convolution, and MLP branches. Each branch independently processes its input using lightweight modules that specialize in distinct inductive biases, i.e., global dependency modeling, local pattern extraction, and channel-wise transformation. This design reduces computation (each operation handles only a fraction of the channels) while enhancing representational diversity through architectural heterogeneity (see Lines 1-8 Alg. 1).
@@
Inter-Branch Mutual Fusion Injects Expressivity Before FFN. While the parallel design improves efficiency and modularity, it inevitably reduces interaction across branches. To mitigate this, Lines 9–13 of Alg. 1 introduce an inter-branch mutual fusion mechanism, where each branch is enhanced via gated aggregation of the other two, modulated by learnable coefficients $\lambda$. This introduces non-linearity and cross-path context blending, reinforcing feature complementarity before unification. The fused representation is more expressive, forming a compact and effective pre-FFN encoder.

### v2

# arXiv:2505.18679v2[cs.CV] 27 Feb 2026

@@ Abstract
Degradation-agnostic image restoration aims to handle diverse corruptions with one unified model, but faces fundamental challenges in balancing efficiency and performance across different degradation types. Existing approaches either sacrifice efficiency for versatility or fail to capture the distinct representational requirements of various degradations. We present MIRAGE, an efficient framework that addresses these challenges through two key innovations. First, we propose a channel-wise functional decomposition that systematically repurposes channel redundancy in attention mechanisms by assigning CNN, attention, and MLP branches to handle local textures, global context, and channel statistics, respectively. This principled decomposition enables degradation-agnostic learning while achieving superior efficiency-performance trade-offs. Second, we introduce manifold regularization that performs cross-layer contrastive alignment in Symmetric Positive Definite (SPD) space, which empirically improves feature consistency and generalization across degradation types. Extensive experiments demonstrate that MIRAGE achieves state-of-the-art performance with remarkable efficiency, outperforming existing methods in various all-in-one IR settings while offering a scalable and generalizable solution for challenging unseen IR scenarios.
@@ 1 Introduction
Image Restoration (IR) aims to recover clean images from inputs degraded by diverse real-world corruptions such as noise, blur, haze, rain, and low-light conditions (Zamir et al., 2022; Li et al., 2023a; Ren et al., 2024; Potlapalli et al., 2024). A central challenge is degradation-agnostic restoration: developing a single model that can generalize across heterogeneous degradations. Despite recent progress, existing approaches often face an efficiency–performance dilemma. On the one hand, heavyweight designs based on prompts, instructions, or large vision–language models provide versatility but incur substantial computational cost (Potlapalli et al., 2024; Zamfir et al., 2025; Jiang et al., 2025). On the other hand, lightweight solutions improve efficiency at the expense of restoration quality (Li et al., 2022; Tang et al., 2025b). Achieving both robustness and efficiency within a unified framework remains an open problem.
This difficulty can be better understood from two complementary perspectives. First, different degradation types impose fundamentally different representational requirements: additive corruptions (e.g., noise, rain) primarily affect local textures, multiplicative distortions (e.g., haze, low-light) require global context modeling, and kernel-based degradations (e.g., blur) call for multi-scale structural reasoning. At the same time, basic architectural modules exhibit distinct inductive biases: convolutional filters excel at local texture modeling, attention mechanisms capture long-range dependencies, and MLPs enhance channel statistics. This motivates the insight that an effective restoration model should systematically align distinct modules with complementary representational functions. Second, recent studies reveal substantial redundancy in attention-based models, particularly along the channel dimension (Venkataramanan et al., 2024; Dong et al., 2021). Many channels encode overlapping information, suggesting that this redundancy could be repurposed rather than discarded. Leveraging this observation allows for architectures that remain compact while preserving expressive capacity. These observations highlight that unified IR benefits not only from adding new modules, but from a principled reorganization of existing capacity based on redundancy patterns and complementary inductive biases. This perspective motivates our design philosophy in MIRAGE, where representational roles are explicitly aligned with structural evidence rather than heuristic module stacking.
Building on these insights, we present MIRAGE, an efficient framework for degradation-agnostic image restoration. MIRAGE introduces two components. (i) Channel-wise functional decomposition, where the input feature map is partitioned along the channel dimension and processed by three specialized branches: convolution for local textures, attention for global context, and MLP for channel statistics. This structured decomposition repurposes redundant capacity into complementary roles, yielding both interpretability and strong efficiency–performance trade-offs. (ii) Manifold regularization, a cross-layer contrastive strategy that leverages natural feature pairs within the model. Inspired by deeply supervised networks (Lee et al., 2015), we hypothesize that natural contrastive pairs exist between shallow and latent representations. Shallow features preserve fine spatial details but are sensitive to noise, while latent features are more abstract and semantically stable; aligning them encourages more robust shared representations. Importantly, rather than computing contrastive loss in Euclidean space, which may distort similarity when comparing structured representations, we operate in the Symmetric Positive Definite (SPD) manifold space. This formulation provides a more faithful alignment of representations, leading to improved generalization across degradation types. Overall, MIRAGE provides a structurally grounded view of unified IR, where representational capacity is allocated and aligned based on statistical evidence at both the spatial and depth levels.
Extensive experiments across five degradation settings show that MIRAGE achieves state-of-the-art performance with remarkable efficiency: our model has only 6M parameters, more than five times smaller than recent prompt-based baselines, while also generalizing well to unseen scenarios such as underwater image enhancement. Both the visual and per-setting PSNR results are shown in Fig. 1.
Our contributions are summarized as follows:
  • We propose a principled channel-wise functional decomposition strategy that aligns convolution, attention, and MLP with distinct representational roles, enabling efficient and effective degradation-agnostic restoration.
  • We introduce manifold regularization through cross-layer contrastive alignment between shallow and latent features. We exploit natural contrastive pairs within the model, and perform this alignment in the SPD manifold space rather than Euclidean space, providing more faithful representation similarity and improved generalization across diverse degradations.
  • We conduct comprehensive experiments across single, mixed, and unseen degradation settings, establishing MIRAGE as a strong and practical baseline for all-in-one IR.
@@ 2 Related Work
Image Restoration with Various Architectures. IR addresses the ill-posed problem of retoring high-quality images from degraded inputs and has long been a core task in computer vision with broad applications (Richardson, 1972; Banham and Katsaggelos, 1997; Xie et al., 2025; Li et al., 2023b; Zamfir et al., 2024). Early methods relied on model-based formulations with handcrafted priors, but deep learning has shifted the field toward data-driven approaches, including regression-based (Lim et al., 2017; Lai et al., 2017; Liang et al., 2021; Chen et al., 2021; Li et al., 2023a; Zhang et al., 2024) and generative pipelines (Gao et al., 2023; Wang et al., 2023b; Luo et al., 2023; Yue et al., 2023; Zhao et al., 2024). These methods employ diverse backbones: convolutional networks for local structures (Dong et al., 2015; Zhang et al., 2017b; a; Wang et al., 2018), MLPs and state space models for channel or sequential dependencies (Tu et al., 2022; Guo et al., 2024a; Zhu et al., 2024; Gu and Dao, 2023; Dao and Gu, 2024; Tang et al., 2025a), and Transformers for long-range interactions (Liang et al., 2021; Ren et al., 2023; Li et al., 2023a; Zamir et al., 2022; Dosovitskiy et al., 2020; Liu et al., 2023; Shi et al., 2025), achieving promising results. Despite these advances, most IR solutions remain degradation-specific, addressing tasks such as denoising (Zhang et al., 2019), dehazing (Wu et al., 2021), deraining (Jiang et al., 2020), or deblurring (Kong et al., 2023), motivating the need for unified frameworks that generalize across diverse degradations while remaining efficient.
@@
Degradation-agnostic Image Restoration. While training task-specific models for individual degradations can be effective, it is impractical to maintain separate models for each corruption. Real-world images often suffer from mixed degradations, making independent treatment infeasible, and task-specific approaches further increase computational and storage costs, amplifying their environmental footprint. To overcome these limitations, the emerging field of degradation-agnostic IR focuses on single-blind models capable of handling multiple degradation types simultaneously (Zamfir et al., 2025; Zeng et al., 2025; Zheng et al., 2024b; Ren et al., 2026). For example, AirNet (Li et al., 2022) achieves blind All-in-One image restoration by using contrastive learning to derive degradation representations from corrupted images, which are then leveraged to reconstruct clean images. Building on this, IDR (Zhang et al., 2023) tackles the problem by decomposing degradations into fundamental physical components and applying a two-stage meta-learning strategy. More recently, the extra learnable prompt-based paradigm (Potlapalli et al., 2024; Wang et al., 2023a; Li et al., 2023c; Tian et al., 2025) has introduced a visual prompt learning module, enabling a single model to better handle diverse degradation types by leveraging the discriminative capacity of learned visual prompts. Extending this idea, some works further model prompts from a frequency perspective (Cui et al., 2025) or propose more complex architectures with additional datasets (Dudhane et al., 2024). However, visual prompt modules often result in increased training time and decreased efficiency (Cui et al., 2025). Meanwhile, inspired by recent advances in self-supervised learning, several works (Wu et al., 2021; Chen et al., 2022c) have explored contrastive objectives to enhance low-level representations, though mainly within single-task IR scenarios. For the degradation-agnostic setting (Jiang et al., 2025; Li et al., 2022; Chen et al., 2025b; Zhang et al., 2025), the most recent DA-RCOT (Tang et al., 2025b) introduces a contrastive loss applied to residual feature space, illustrating that contrastive signals can also benefit unified IR models. In contrast, our work aims to improve the model’s ability to capture representative degradation cues within the SPD space without relying on heavy or complex prompt designs. Our goal in this work is to develop a degradation-agnostic image restorer that remains both computationally efficient and environmentally sustainable.
3 Preliminary: Degradation-Aware Architectures for IR
@@
Image Degradation and Restoration. Image restoration seeks to recover a clean image $\mathbf{x}$ from a degraded observation $\mathbf{y}$:
@@
where $\mathcal{D}(\cdot)$ denotes a degradation operator and $\mathbf{n}$ noise. Real-world degradations are diverse—additive (e.g., Gaussian noise, rain: $\mathbf{y}=\mathbf{x}+\mathbf{n}$), multiplicative (e.g., haze, speckle: $\mathbf{y}=\mathbf{x}\cdot\mathbf{m}$), or convolutional (e.g., blur, super-resolution: $\mathbf{y}=\mathbf{k}*\mathbf{x}+\mathbf{n}$(He et al., 2025). These factors often co-occur and are spatially variant (Zhai et al., 2023), forming compound pipelines:
$\mathbf{y}=\mathcal{D}_{3}\big(\mathcal{D}_{2}(\mathcal{D}_{1}(\mathbf{x}))\big)+\mathbf{n}.$ (2)
Such complexity demands models that preserve local details while reasoning about global structures.
Architectural Biases for Degradation Modeling. Deep networks embody distinct inductive biases: CNNs capture local spatial patterns: $\mathbf{y}p=\sum{i\in\mathcal{N}(p)}w_{i}\cdot\mathbf{x}_{i}$, effective for uniform or spatially invariant degradations. Transformers exploit global self-attention: $\mathbf{y}i=\sum_{j}\alpha{ij}\cdot\mathbf{V}_{j}$, well-suited for non-uniform, structured degradations (e.g. haze, patterned noise). MLPs, especially token-mixing forms, apply flexible position-wise mappings: $\mathbf{y}=\mathbf{W}_{2}\cdot\phi(\mathbf{W}_{1}\cdot\mathbf{x})$, though with weak spatial priors.
Each paradigm shows strengths yet clear limitations—CNNs excel in local fidelity, Transformers in global reasoning, and MLPs in flexible feature interactions, but lack inductive structure. Alone, they are insufficient for complex degradations and often parameter-heavy. Their complementarity motivates unified, degradation-aware architectures that leverage all three for robust IR in the wild.
@@ 4 The Proposed MIRAGE
Figure 2: (a)-(c): The most adopted all-in-one image restoration encoder-decoder pipelines. (d): The toy illustration of our SPD contrastive pipeline. (e): The overall framework of the proposed MIRAGE : i.e., a convolutional patch embedding, a U-shape encoder-decoder main body, an extra refined block, and the proposed SPD contrastive learning algorithm. (f): Structure of each mixed degradation adaptation block (MDAB).
The design of MIRAGE is guided by two empirical observations. (i) Attention features consistently exhibit low-rank channel redundancy across scales (Fig. 3), indicating that a non-trivial portion of the representational capacity can be reassigned without loss of expressiveness. (ii) Different degradations favor complementary inductive biases, i.e., local texture sensitivity, global contextual aggregation, and channel-statistical modulation. These observations motivate a principled partition of feature channels into convolutional, attention, and MLP pathways, allowing each subspace to specialize in the bias it is best suited for while maintaining overall model compactness. In parallel, the depth-asymmetric covariance structures of shallow and latent representations provide a natural basis for cross-layer alignment, for which the SPD formulation offers a geometry-preserving representation.
Prior works either train a separate model per degradation (Fig. 2a), adopt multi-encoder–single-decoder designs that inflate parameters (Fig. 2b), or rely on large-scale prompt-based models with visual/textual cues (Fig. 2c). In contrast, we propose a simple yet effective mixed-backbone architecture (Fig. 2d), which already forms a strong restoration baseline (Sec. 4.1) and is further enhanced by cross-layer contrastive learning in SPD space between shallow and latent features (Sec. 4.2).
4.1 Mixed Degradation Adaptation Block for Degradation-Agnostic IR
@@
Redundancy in MHAs Opens Opportunities for Hybrid Architectures. Redundancy has long been recognized as a fundamental limitation in multi-head self-attention (MHA), the core building block of Transformers, in both NLP and vision domains (Nguyen et al., 2022b; a; Xiao et al., 2024; Brödermann et al., 2025; Wang et al., 2022; Venkataramanan et al., 2024). Prior studies indicated that not all attention heads contribute equally, i.e., some are specialized and crucial, while others can be pruned with negligible impact. This inherently implies redundancy in the channel dimension, as MHA outputs are concatenated along this axis. To empirically verify this redundancy in the context of IR, we analyze intermediate features from a lightweight attention-only model (details in the Appendix A). Specifically, we compute cumulative explained variance via PCA and normalized singular value spectra via SVD across multiple feature scales. Fig. 3(a) shows earlier scales (e.g., 1st Scale) need far fewer principal components to retain most variance, suggesting high redundancy. Fig. 3(b) further supports this, with a sharper singular value decay at shallower stages, indicating stronger low-rank structure in channel-wise representations. Even at the deepest stage (e.g., 4th Scale), achieving 90% variance requires only 31 of 192 components ($\approx 16\%$), confirming redundancy persists throughout.
This insight motivates a departure from traditional head/channel pruning. Instead of discarding redundant capacity, we propose to repurpose it by splitting the channel dimension into three parts and feeding them into distinct architectural branches, i.e., attention, convolution, and MLP. This hybrid formulation leverages complementary inductive biases and makes full use of available representational space, offering a principled and efficient alternative to the previous pure MSA-based designs.
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
Parallel Design Brings More Efficiency. As shown in Lines 1–8 of Alg. 1, we instantiate this idea through a structurally parallel design that simultaneously exploits complementary inductive biases. As illustrated in Fig. 2(f), the input feature $F_{\text{in}}\in\mathbb{R}^{h\times w\times c}$ is evenly partitioned along the channel dimension into three sub-tensors (i.e., $F_{\text{in}}^{\text{att}}$, $F_{\text{in}}^{\text{conv}}$, and $F_{\text{in}}^{\text{mlp}}$,), which are then processed in parallel by attention, convolution, and MLP branches. Each branch operates only on its allocated fraction of channels, substantially reducing computational cost, while its architectural heterogeneity enriches the representational space. This parallel decomposition achieves a favorable balance between efficiency and expressiveness, in contrast to prior designs that rely on purely attention-based processing.
@@
Inter-Branch Mutual Fusion Injects Expressivity Before FFN. While the parallel design improves efficiency and modularity, it reduces interaction across branches. To mitigate this, Lines 9–13 of Alg. 1 introduce an inter-branch fusion mechanism, where each branch is enhanced via gated aggregation of the rest, modulated by learnable coefficients $\lambda$. This introduces cross-path context blending, reinforcing feature complementarity before unification, forming an effective pre-FFN encoder.